OpenStudy (anonymous):

Decide whether the given ordered pair is a solution of the given system. X - Y= 17 ; (8,9) X + Y=-1

OpenStudy (igreen):

Plug in the values and see if they satisfy both equations.

OpenStudy (igreen):

x - y = 17 x + y = -1 (x, y) 8 - 9 = 17 8 + 9 = -1 See if those equations are true.

OpenStudy (igreen):

Just simplify the left side of the equal sign and see if it is equal to the right side of the equal sign.